Please HELP me!! I am TOTALLY confused! Here is the problem♥:
Franklin can do a job in 6 h and Mike can do it in 4 h. What part of the job can they do by working together for 2 h? for y h?
♥Thanx sOo much♥

Answer by venugopalramana(3286)


Franklin can do a job in 6 h
HENCE FRANKLIN CAN DO IN 1 HOUR .......1/6 JOB
and Mike can do it in 4 h.
HENCE MIKE CAN DO IN 1 HOUR...........1/4 JOB
HENCE TOGETHER THEY CAN DO IN 1 HOUR ....1/6 + 1/4 = (2+3)/12=5/12 JOB
What part of the job can they do by working together for 2 h?
HENCE IN 2 HOURS THEY CAN DO TOGETHER........2*5/12=5/6 JOB...
for y h?
HENCE IN Y HOURS THEY CAN DO TOGETHER........Y*5/12=5Y/6 JOB...